The mean and variance of a Binomial variable X with parameters n and p are 16 and 8. Find P(X1) and P(X>2).

Answer by Boreal(15235)

the mean is np=16
the variance is np (1-p)=8
so 16(1-p)=8
1-p=1/2
p=0.50
n=32
p=0.5
mean 16
variance 8
probability x=1 is 32C)*0.5^1*0.5^31=7 x 10^(-9)
probability x>2 is 1- probability x=2 essentially, which is 1=32C2*0.5^32=1.